the drosophila chromosome #3 is 27.9 million base pairs long.
1. how many nucleosomes and histones would be present on one of these chromosomes at the mitotic metaphase plate?
2. if this chromosome is 30% G/C, how many hydrogen bonds and phosphodiester linkages would a single chromosome have before the S phase?
3. how many of these G0 chromosomes would you need to stretch end to end to equal one kilometer?
Answer #1

1). 1 nucleosome = 146 bp

So nucleosomes in that chromosome= 27.9 million / 146

ie 191095.89 nucleosomes

1 nucleosome consists of 8 histone proteins. So histones in that chromosome = 1528767 histone proteins

2) 1 GC linkage has 3 hydrogen bonds. 30% GC means 27900000 X .3 = 8370000 bp connected by 3 hydrogen bonds nd rest by 2 hydrogen bonds..so total hydrogen bonds 8370000 X 3 = 25110000 + 19,530,000 X 2

So total = 64170000 hydrogen bonds

In DNA and RNA, the phosphodiester bond is the linkage between the 3' carbon atom of one sugar molecule and the 5' carbon atom of another, deoxyribose in DNA and ribose in RNA.

No of phosphodiester bonds = total nucleotides -2

Nucleotides in above chromosome = 2 X no of base pairs

27900000*2 = 55800000

So phosphodiester linkage= 55800000-2 55799998

3)о 3 Ibp= 3.4 À 3.4x100 m. So chromosomes reg for 1km be required for 11 km = & 1000ml 100x10 - 2.94 x 102 bp 1000 3.4 1 Chro